OpenStudy (keana):

simplify: √2(x-√5) a: 10 √2x b: 2x√2 c: √2x-√10 d: x√2-√10

OpenStudy (anonymous):

Distribute the radical to both the inner terms and tell me what you got??

OpenStudy (keana):

@waterineyes though i don't know how !

OpenStudy (anonymous):

See, \[\sqrt{a}(x - \sqrt{b}) = \sqrt{a}.x - \sqrt{a}.\sqrt{b}\] Here, a = 2 and b = 5 Use this formula...

OpenStudy (anonymous):

Have you tried the above formula or you are only looking for answers??

OpenStudy (keana):

why are there periods after the a's? @waterineyes

OpenStudy (anonymous):

No tell me why are you only bumping the question again and again instead of trying it???

OpenStudy (keana):

@waterineyes you mean once .what you said didn't work & idk why there are periods .

OpenStudy (anonymous):

Did not work??? Show me your work and then I tell you where you have mistaken.. And I am helping others too...

OpenStudy (anonymous):

Now do you want to solve or just looking that someone will come and solve it for you??

OpenStudy (keana):

i am trying to answer the problem so don't rush me k

OpenStudy (keana):

the answer i get doesn't work

OpenStudy (anonymous):

Sure take your time...

OpenStudy (anonymous):

But what you get???

OpenStudy (keana):

\[2√5+25=2x\]

OpenStudy (keana):

this is what i get

OpenStudy (anonymous):

Yes but you have to remember formula @keana How much time has been past since I told you the formula that: \[\sqrt{a}.\sqrt{b} = \sqrt{a \times b}\] You get: \[\sqrt{2}.\sqrt{5}\] Can you now solve and don't forget the formula after solving the problem please..

OpenStudy (anonymous):

In the attachment you have done the right.. Use the above formula to get the answer...

OpenStudy (keana):

multiplying ?

OpenStudy (anonymous):

What? Tell me the answer you are getting in the attachment you have attached.. Just tell that and forget about others...

OpenStudy (keana):

am i multiplying is what i'm asking

OpenStudy (anonymous):

You got right till here.. \[\sqrt{2}.x - \sqrt{2}.\sqrt{5}\] You have to solve this further.. See use the above formula that is: \[\sqrt{a}.\sqrt{b} = \sqrt{a \times b}\] Just use this formula in the answer you get..

OpenStudy (keana):

10 ?

OpenStudy (anonymous):

Are you sure? Why you always forget to say it as: \[\sqrt{10}\] ???

OpenStudy (keana):

lazy typing ?

OpenStudy (anonymous):

You are right now what option did you get??

OpenStudy (keana):

what do you mean by option ?

OpenStudy (anonymous):

What is the answer now from the choices given???

OpenStudy (keana):

a ?

OpenStudy (anonymous):

How dear??

OpenStudy (keana):

i don't know what the answer is ! that's why i asked !!

OpenStudy (anonymous):

Now tell me finally who got as answer?? \[\sqrt{2}.x - \sqrt{2}.\sqrt{5} = \sqrt{2}.x - \sqrt{10}\] Please do me a favour now can you please tell me is it a, b, c or d??

OpenStudy (anonymous):

Just reply what are you waiting for??

OpenStudy (keana):

i want to see if i can get it without guessing

OpenStudy (anonymous):

See =, options in more understandable form: c option is like this: \[\sqrt{2x} - \sqrt{10}\] d is like this: \[x \sqrt{2} - \sqrt{10}\]

OpenStudy (keana):

c ?

OpenStudy (anonymous):

is x in root??

OpenStudy (anonymous):

Well one thing I want to say that you don't how to solve your previous questions and you are bumping more questions...

OpenStudy (anonymous):

The answer is D.. And thanks for your cooperation...